# TTOM: Test-Time Optimization and Memorization for Compositional Video Generation

### v1

@@ Abstract
Video Foundation Models (VFMs) exhibit remarkable visual generation performance, but struggle in compositional scenarios (e.g., motion, numeracy, and spatial relation). In this work, we introduce Test-Time Optimization and Memorization (TTOM), a training-free framework that aligns VFM outputs with spatiotemporal layouts during inference for better text-image alignment. Rather than direct intervention to latents or attention per-sample in existing work, we integrate and optimize new parameters guided by a general layout-attention objective. Furthermore, we formulate video generation within a streaming setting, and maintain historical optimization contexts with a parametric memory mechanism that supports flexible operations, such as insert, read, update, and delete. Notably, we found that TTOM disentangles compositional world knowledge, showing powerful transferability and generalization. Experimental results on the T2V-CompBench and Vbench benchmarks establish TTOM as an effective, practical, scalable, and efficient framework to achieve cross-modal alignment for compositional video generation on the fly.
@@ 1 Introduction
Recent years have witnessed a rapid stride in video generation (Ho et al., 2022; Polyak et al., 2024) and world simulation (Brooks et al., 2024). Benefiting from the significant advance in flow matching (Lipman et al., 2022; Esser et al., 2024) and diffusion transformers (DiTs) (Peebles & Xie, 2023), current text-to-video generation (T2V) models (Yang et al., 2024b; Kong et al., 2024) are able to synthesize realistic and vivid videos. However, even state-of-the-art models still suffer from text-video misalignment in compositional scenarios, i.e., composing multiple objects, attributes, and relations into a complex scene (Sun et al., 2025; Huang et al., 2024).
Remarkable efforts have been made to improve text-video alignment by introducing explicit spatiotemporal layout guidance (Lian et al., 2023b; He et al., 2025) as a bridge between text prompts and videos. In this paradigm, In this paradigm, a layout represented by bounding box (bbox) sequences—each corresponding to an object—is first generated by a large language model (LLM), and subsequently used to optimize intermediate representations (e.g., latents and attention maps) by enforcing alignment with the layout constraints. Despite notable progress, this line of work faces several limitations: 1) direct intervention on intermediate representations may disrupt feature distributions, leading to degraded video quality (e.g., inconsistency, flickering artifacts, or even collapse); 2) the per-sample control paradigm treats each case independently, thereby neglecting the historical context of previously generated videos; and 3) these methods do not enhance the intrinsic capability of generative models, since interventions on one sample fail to generalize to others.
@@ 2 Related Work
Compositional Visual Generation. Despite the thrilling progress in foundation models (Yang et al., 2024b; Kong et al., 2024; Polyak et al., 2024; Seawead et al., 2025), compositional generation (Chefer et al., 2023; Feng et al., 2022) remains an open challenge and has attracted substantial research interest. To enhance text-image alignment in compositional scenes, a rich line of studies (Qu et al., 2023; Feng et al., 2023) explores incentivizing the layout planning ability of LLMs (Brown et al., 2020; Ouyang et al., 2022) to enable layout-guided controllable generation (Tian et al., 2024; Lin et al., 2023; Feng et al., 2025) through attention control (Chen et al., 2024; Xie et al., 2023; He et al., 2023; Wang et al., 2024; He et al., 2025) or latent modification (Qi et al., 2024; Wu et al., 2024; Lian et al., 2023a; Yang et al., 2024a; Wang et al., 2025b; Lian et al., 2023b), typically in a training-free manner. However, direct intervention on attention maps or latents may compromise visual quality. Moreover, the per-sample paradigm treats test samples in isolation, disregarding informative historical context.
@@
Memory Mechanism. To track and model historical contexts, existing Test-Time Training (TTT) (Sun et al., 2024) work designs memory mechanisms instantiated with inserted layers, i.e., TTT layers. The memory can be updated linearly (Sun et al., 2020) or in a local temporal window (Wang et al., 2025a), by optimizing TTT layers via a self-supervised learning loss (He et al., 2022). Despite the remarkable advance, they often consider intra-sample context (e.g., frames in video) (Wang et al., 2025a; Dalal et al., 2025) and assume the independence among samples. Besides, the memory has a fixed size and does not support flexible operations.
@@ 3.1 LLM-driven Spatial-Temporal Layout Planning
Text-to-Layout Generation. Prior work has shown that LLMs possess strong spatial-temporal dynamics understanding abilities (Qu et al., 2023; Feng et al., 2023), and capture physical properties (Lian et al., 2023b), such as gravity, elasticity, and perspective projection, which align closely with real-wrold physics. Inspired by it, we prompt LLMs to generate SPLs given user prompts via in-context learning (Dong et al., 2022). To enhance scene comprehension, the LLM is first instructed to produce descriptions of object motions and camera behaviors, followed by the corresponding layout generation. Finally, a verification step ensures spatial and temporal consistency of object phrases and bounding box sequences, with corrections applied when discrepancies are detected.
@@ 3.2 Test-Time Optimization for Layout-to-Video Generation
In this section, we first discuss the relation between intermediate attention maps of diffusion models and generated videos, and conduct a probe experiment to verify the relevance. On the basis of the relevance, we propose a training-free controllable generation strategy by optimizing attention-to-layout alignment during inference, i.e., Test-Time Optimization (TTO).
Attention-Layout Relevance Probe in Diffusion Transformers. Prior work (Hertz et al., 2022) has demonstrated the relevance between cross-attention maps and spatial layouts, which is exploited for video editing (Qi et al., 2023) and controllable generation (Xie et al., 2023). However, with the architecture shift from UNets (Ronneberger et al., 2015) to Diffusion Transformers (DiTs) (Peebles & Xie, 2023) due to its stronger scalability, the recent study (Wang et al., 2025b) pointed out their difference, i.e., the cross-attention maps in DiTs distribute more evenly than those in UNets, which may cause the ineffectiveness of previous methods due to the weak attention-layout relevance.
Motivated by it, we propose a probe strategy to assess attention-layout relevance across different layers: 1) preparing a set of prompts describing a common scene in which an object performs a motion. 2) generating videos given the prompts, and meanwhile, extracting text-video cross-attention maps. Specifically, we first extract the cross-attention map††For foundation models (Yang et al., 2024b) without cross-attention layers, we extract those parts in self-attention maps where text-vision interaction happens. from each layer for the object phrase from a text prompt , where denotes the token number of . , , and refer to the frame number, height, and width of the video latent, respectively. Afterwards, we perform average pooling along the text dimension and obtain . This process is formulated as:
@@ 3.3 Parametric Memory
Discussion. Compared with prior methods (Lian et al., 2023b; He et al., 2025) treating each sample independently, the advantages of the proposed memory mechanism lie in: 1) Superiority. The history optimization maintained in the memory provides abundant scene knowledge, serving as good initialization for TTO. 2) Personalization. In practice, we can maintain a user-specific memory for each user, track their historical requests, and model the intention for better personalized generation. 3) Efficiency. For those prompts similar to historical ones, we can skip optimization and directly read and load items into the foundation model for inference. 4) Scalability. We can scale the memory up in two dimensions: increasing the item-level capacity (i.e., enlarging the history context window) could memorize more historical optimization information; and increasing the number of per-item parameters may capture more scene patterns in each TTO.
@@ 4.1 Experimental Setup
Streaming Video Generation Setting. To align with the practical video creation scenarios in which streaming user prompts come one by one, we propose to model historical streaming contexts. To strictly keep a fair comparison with existing methods that view test samples independently, we have a test-time independence setting: we first generate 200 user prompts using GPT-4o to cover common compositional scenes, and then generate videos with these prompts, at the same time, perform TTOM to construct a memory. During inference, we can read and load entries from the memory, but forbid inserting or updating the memory with the optimization results of test samples.

### v2

@@ Abstract
Video Foundation Models (VFMs) exhibit remarkable visual generation performance, but struggle in compositional scenarios (e.g., motion, numeracy, and spatial relation). In this work, we introduce Test-Time Optimization and Memorization (TTOM), a model-agnostic framework that aligns VFM outputs with spatiotemporal layouts during inference for better text-image alignment. Rather than direct intervention to latents or attention per-sample in existing work, we integrate and optimize new parameters guided by a general layout-attention objective. Furthermore, we formulate video generation within a streaming setting, and maintain historical optimization contexts with a parametric memory mechanism that supports flexible operations, such as insert, read, update, and delete. Notably, we found that TTOM disentangles compositional world knowledge, showing powerful transferability and generalization. Experimental results on the T2V-CompBench and Vbench benchmarks establish TTOM as an effective, practical, scalable, and efficient framework to achieve cross-modal alignment for compositional video generation on the fly.
@@ 1 Introduction
Recent years have witnessed a rapid stride in video generation (Ho et al., 2022; Polyak et al., 2024) and world simulation (Brooks et al., 2024). Benefiting from the significant advance in flow matching (Lipman et al., 2022; Esser et al., 2024) and diffusion transformers (DiTs) (Peebles and Xie, 2023), current text-to-video generation (T2V) models (Yang et al., 2024b; Kong et al., 2024) are able to synthesize realistic and vivid videos. However, even state-of-the-art models still suffer from text-video misalignment in compositional scenarios, i.e., composing multiple objects, attributes, and relations into a complex scene (Sun et al., 2025; Huang et al., 2024).
Remarkable efforts have been made to improve text-video alignment by introducing explicit spatiotemporal layout guidance (Lian et al., 2023b; He et al., 2025) as a bridge between text prompts and videos. In this paradigm, a layout represented by bounding box (bbox) sequences—each corresponding to an object—is first generated by a large language model (LLM), and subsequently used to optimize intermediate representations (e.g., latents and attention maps) by enforcing alignment with the layout constraints. Despite notable progress, this line of work faces several limitations: 1) direct intervention on intermediate representations may disrupt feature distributions, leading to degraded video quality (e.g., inconsistency, flickering artifacts, or even collapse); 2) the per-sample control paradigm treats each case independently, thereby neglecting the historical context of previously generated videos; and 3) these methods do not enhance the intrinsic capability of generative models, since interventions on one sample fail to generalize to others.
@@ 2 Related Work
Compositional Visual Generation. Despite the thrilling progress in foundation models (Yang et al., 2024b; Kong et al., 2024; Polyak et al., 2024; Seawead et al., 2025), compositional generation (Chefer et al., 2023; Feng et al., 2022; Qu et al., 2024b; 2025b) remains an open challenge and has attracted substantial research interest. To enhance text-image alignment in compositional scenes, a rich line of studies (Qu et al., 2023; Feng et al., 2023) explores incentivizing the layout planning ability of LLMs (Brown et al., 2020; Ouyang et al., 2022) to enable layout-guided controllable generation (Tian et al., 2024; Lin et al., 2023; Feng et al., 2025) through attention control (Chen et al., 2024; Xie et al., 2023; He et al., 2023; Wang et al., 2024; He et al., 2025) or latent modification (Qi et al., 2024; Wu et al., 2024; Lian et al., 2023a; Yang et al., 2024a; Wang et al., 2025c; Lian et al., 2023b), typically in a training-free manner. However, direct intervention on attention maps or latents may compromise visual quality. Moreover, the per-sample paradigm treats test samples in isolation, disregarding informative historical context.
@@
Memory Mechanism. To track and model historical contexts, existing Test-Time Training (TTT) (Sun et al., 2024) work designs memory mechanisms instantiated with inserted layers, i.e., TTT layers. The memory can be updated linearly (Sun et al., 2020) or in a local temporal window (Wang et al., 2025b), by optimizing TTT layers via a self-supervised learning loss (He et al., 2022). Despite the remarkable advance, they often consider intra-sample context (e.g., frames in video) (Wang et al., 2025b; Dalal et al., 2025) and assume the independence among samples. Besides, the memory has a fixed size and does not support flexible operations.
@@ 3.1 LLM-driven Spatial-Temporal Layout Planning
Text-to-Layout Generation. Prior work has shown that LLMs possess strong spatial-temporal dynamics understanding abilities (Qu et al., 2021; 2023; Feng et al., 2023; Qu et al., 2024a), and capture physical properties (Lian et al., 2023b), such as gravity, elasticity, and perspective projection, which align closely with real-world physics. Inspired by it, we prompt LLMs to generate SPLs given user prompts via in-context learning (Dong et al., 2022; Qu et al., 2025a). To enhance scene comprehension, the LLM is first instructed to produce descriptions of object motions and camera behaviors, followed by the corresponding layout generation. Finally, a verification step ensures spatial and temporal consistency of object phrases and bounding box sequences, with corrections applied when discrepancies are detected.
@@ 3.2 Test-Time Optimization for Layout-to-Video Generation
In this section, we first discuss the relation between intermediate attention maps of diffusion models and generated videos, and conduct a probe experiment to verify the relevance. On the basis of the relevance, we propose a model-agnostic controllable generation strategy by optimizing attention-to-layout alignment during inference, i.e., Test-Time Optimization (TTO).
Attention-Layout Relevance Probe in Diffusion Transformers. Prior work (Hertz et al., 2022) has demonstrated the relevance between cross-attention maps and spatial layouts, which is exploited for video editing (Qi et al., 2023) and controllable generation (Xie et al., 2023). However, with the architecture shift from UNets (Ronneberger et al., 2015) to Diffusion Transformers (DiTs) (Peebles and Xie, 2023) due to its stronger scalability, the recent study (Wang et al., 2025c) pointed out their difference, i.e., the cross-attention maps in DiTs distribute more evenly than those in UNets, which may cause the ineffectiveness of previous methods due to the weak attention-layout relevance.
Motivated by it, we propose a probe strategy to assess attention-layout relevance across different layers: 1) preparing a set of prompts describing a common scene in which an object performs a motion. 2) generating videos given the prompts, and meanwhile, extracting text-video cross-attention maps. Specifically, we first extract the cross-attention map111For foundation models (Yang et al., 2024b) without cross-attention layers, we extract those parts in self-attention maps where text-vision interaction happens. from each layer for the object phrase from a text prompt , where denotes the token number of . , , and refer to the frame number, height, and width of the video latent, respectively. Afterwards, we perform average pooling along the text dimension and obtain . This process is formulated as:
@@ 3.3 Parametric Memory
Discussion. Compared with prior methods (Lian et al., 2023b; He et al., 2025) treating each sample independently, the advantages of the proposed memory mechanism lie in: 1) Superiority. The history optimization maintained in the memory provides abundant scene knowledge, serving as good initialization for TTO. 2) Personalization. In practice, we can maintain a user-specific memory for each user, track their historical requests, and model the intention for better personalized generation (Wang et al., 2025a; Zhao et al., 2025; 2026). 3) Efficiency. For those prompts similar to historical ones, we can skip optimization and directly read and load items into the foundation model for inference. 4) Scalability. We can scale the memory up in two dimensions: increasing the item-level capacity (i.e., enlarging the history context window) could memorize more historical optimization information; and increasing the number of per-item parameters may capture more scene patterns in each TTO.
@@ 5 Conclusion
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. A link (https://ttom-t2v.github.io/) to the source code is provided, enabling replication of our implementation. The main text and appendix together provide comprehensive descriptions of the model design, optimization procedure, and evaluation protocol. These resources collectively ensure that readers can reproduce and validate our experimental results.
@@ B.2 Additional Qualitative Results
Appendix C Prompt Design for LLM Planning
This appendix details the system prompts used in the first stage (LLM Planning) of our pipeline. The planning process involves prompt enrichment, metadata extraction, layout generation, and a dedicated verification stage to ensure consistency.
C.1 Prompt Enrichment
The following system prompt is used to expand short user inputs into detailed video descriptions.
C.2 Metadata Extraction
This prompt extracts structured object information from the enriched description.
C.3 Layout Generation
This prompt generates bounding box sequences based on the prompt and metadata.
C.4 Verification
To ensure the robustness of the generated plan, we employ two distinct verification steps: one for textual metadata and one for spatiotemporal layout.
C.4.1 Metadata Verification Prompt
This step corrects format errors and ensures linguistic guidelines are followed.
C.4.2 Layout Verification Prompt
This step critiques the generated bounding boxes, focusing on **numerical accuracy** and **spatiotemporal consistency**.
Appendix D Theoretical Analysis of the Memory Mechanism
In this section, we provide a formal theoretical analysis of the proposed memory mechanism. We frame the mechanism not merely as a caching strategy, but as a Manifold-Aware Initialization scheme that accelerates convergence by leveraging the local smoothness of the task-parameter mapping. We further analyze the stability of the update rule to demonstrate its robustness against catastrophic forgetting.
D.1 Problem Formulation: TTO as Trajectory Optimization
Let denote the manifold of compositional video generation tasks (user prompts), and let represent the parameter space of the trainable modules (e.g., LoRA weights). We define a mapping such that for any task , there exists an optimal parameter set that minimizes the alignment objective :
The standard Test-Time Optimization (TTO) process seeks to find via Gradient Descent starting from a generic initialization . The memory is defined as a discrete set of tuples , where is the semantic representation (key) of a historical task, and is the converged parameter state (value).
D.2 Convergence Analysis: The Initialization Advantage
The efficiency of our method stems from the bounding of the optimization trajectory. We rely on the Task-Parameter Smoothness Assumption: we assume that if two tasks and are semantically close in the embedding space (i.e., ), their corresponding optimal parameters lie within a local neighborhood on the parameter manifold: .
Under the assumption that the loss landscape is locally -strongly convex and -smooth around the optimum (a standard assumption in fine-tuning analysis), the convergence rate of Gradient Descent at step is bounded by the initial distance to the solution:
where is the learning rate.
In our proposed mechanism, the retrieval (Liu et al., 2018; Wen et al., 2024) function selects an initialization such that:
When a similar structural layout is retrieved, . Consequently, the number of steps required to reach an error tolerance is significantly reduced. This theoretical bound formally explains the reduction in inference latency and the enhanced motion scores observed in our empirical results.
D.3 Stability and Orthogonal Information Retention
A critical challenge in online learning is Catastrophic Forgetting, where updates for a new task degrade performance on previous tasks. We prove that our mechanism is immune to this phenomenon via its discrete storage topology.
Unlike continual learning approaches that update a shared parameter set , our memory update mechanism is additive and orthogonal. The update rule is formally defined as:
Because each optimized parameter set is stored as a discrete entry effectively isolated in the memory bank, the gradient updates for task do not alter the values associated with historical keys .
Forgetting Dynamics.
The only information loss in the system is governed by the cache replacement policy (e.g., Least Frequently Used) which occurs only when the memory capacity exceeds the maximum limit . This ensures that the system prioritizes the retention of high-probability compositional structures, effectively approximating the optimal storage policy for the distribution of user prompts.